OpenStudy (anonymous):

Find 3x - y - 3z if x = -2, y = 1, and z = -2

OpenStudy (anonymous):

3(-2)-1-3(-2)= -6-1+6= -1

OpenStudy (anonymous):

the answer is -1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks